OpenStudy (anonymous):

what is the answer, 10/2x^2

OpenStudy (anonymous):

5x^2

OpenStudy (anonymous):

what is the question?

OpenStudy (shamil98):

Yeah, what is the question? find the derivative? solve for x? simplify?..

OpenStudy (anonymous):

are you supposed to reduce the fraction \[\frac{10}{2x^2}\]?

OpenStudy (anonymous):

yes, reduce the fraction.

OpenStudy (anonymous):

what is \(\frac{10}{2}\) ?

OpenStudy (anonymous):

so 5x^2 right guys?

OpenStudy (shamil98):

10/2x^2 what is 10/2? that would be your numerator.

OpenStudy (shamil98):

No.

OpenStudy (anonymous):

it isn't like 10 over 2x^2 ?

OpenStudy (anonymous):

the \(x^2\) part does not migrate to the numerator, it stays in the denominator

OpenStudy (shamil98):

what satellite said.^

OpenStudy (anonymous):

so.. what the final answer?

OpenStudy (shamil98):

10/2x^2 the x^2 stays in the demoninator you can get rid of the 2 by 10/2

OpenStudy (anonymous):

so the answer is, 5/x^2?

OpenStudy (anonymous):

Thats what I wrote.....

OpenStudy (anonymous):

the numerator is 5 and the denominator is x^2, correct?

OpenStudy (anonymous):

thanks guys..

OpenStudy (shamil98):

yes your final answer is 5/x^2

OpenStudy (anonymous):

Wow.... that is the first thing I wrote!

OpenStudy (shamil98):

no, you put 5x^2

OpenStudy (shamil98):

\[\frac{ 5 }{ x^2 }\]

OpenStudy (shamil98):

is the answer

OpenStudy (anonymous):

Forgot the slash....

OpenStudy (shamil98):

not \[5x^2\]

OpenStudy (anonymous):

I meant to put a slash :/